### v1

@@ 1 $P\leftarrow\text{generate\_prompt}(Y)$
3       if ``$person"\in Y$ then
[…]
@@
[…]
@@ 8       $score_{df}\leftarrow\text{classifier\_score}(X_{df})$
[…]
@@ 3.4 Final Training Dataset of WSSS
After selecting the high-quality generated training samples, the synthetic dataset $\mathcal{D}_{aug}$ and the original dataset $\mathcal{D}_{origin}$ are combined as an extended dataset $\mathcal{D}_{final}$ for the training of WSSS: $\mathcal{D}_{final}=\mathcal{D}_{origin}\cup\mathcal{D}_{aug}$. .
@@ 4 Experiments
In this section, we first present the details of the dataset, evaluation metrics, and implementation. Next, we compare our IAA with state-of-the-art methods on the PASCAL VOC 2012 and MS COCO 2014 benchmarks. Finally, we conduct ablation studies to demonstrate the effectiveness of the proposed method.
Dataset and Evaluated Metric. Our experiments are conducted on the PASCAL VOC 2012 dataset [8], which includes $21$ categories, and MS COCO 2014 [16], featuring $81$ classes. The PASCAL VOC 2012 dataset is commonly expanded using the SBD dataset [10]. For training on PASCAL VOC 2012, we use 10,582 images with image-level labels and 1,449 images for the validation set. In the case of the MS COCO 2014 dataset, approximately 82k images are used for training and around 40k images for validation, with the training images having only image-level annotations. The mean Intersection-over-Union (mIoU) is used as the evaluation metric.
@@
Implementation Details. Our IAA integrates knowledge from pre-trained GPT-4o, Stable Diffusion [21], and ControlNet [33]. We utilize ViT-B/16 as the ViT model. To facilitate the training of the patch-based image classifier, images are resized to 384×384 as [14], and the 24×24 encoded patch features are preserved as input. The model is trained for up to 80 epochs with a batch size of 16, using $\epsilon=0.9$ as the high-quality threshold for both text and image. Our final training dataset serves as input for the WSSS framework, while keeping all other settings consistent with ViT-PCM [22]. The experiments were conducted using two NVIDIA 4090 GPUs. Finally, we used the same verification tasks and settings as ViT-PCM [22].
4.1 Final Segmentation Performance
Table 1: Semantic Segmentation Performance Comparison (mIoU) on MS COCO 2014 and PASCAL VOC 2012.
Model Pub. Backbone VOC COCO
TSCD [32] AAAI23 MiT-B1 67.3 40.1
SAS [13] AAAI23 ViT-B/16 69.5 44.5
FPR [4] ICCV23 ResNet38 70.0 43.9
ToCo [24] CVPR23 ViT-B 70.2 42.3
SFC [35] AAAI24 ViT-B/16 71.2 44.6
IACD [29] ICASSP24 ViT-B/16 71.4 44.3
PGSD [9] TCSVT24 ViT-B/16 68.7 43.9
IAA Ours ViT-B/16 72.3 45.3
Comparison with State-of-the-arts. To evaluate our approach, we train fully supervised models using the pseudo labels from our generated images and compare the segmentation performance with other SOTA methods. As shown in Tab. 1, our method outperforms existing SOTA approaches in final segmentation performance.
Visualization of results. Our IAA effectively leverages an augmentation agent, integrating GPT-4o and diffusion knowledge, to enrich the training data for WSSS. Fig. 5 presents the augmented data used as additional training input. Furthermore, Fig. 4 showcases visualization examples of segmentation outputs. It can be observed that our method consistently achieves more precise segmentation results.
@@ 4.2 Ablation Studies
Table 2: Ablation studies on main components of the proposed framework. DA: Diffusion Augmentation; ISR: Image Self-Refinement; PSR: Prompt Self-Refinement on the Pascal VOC 2012 val.
@@
We conduct ablation studies to assess the impact of our two key contributions: prompt self-refinement and image self-refinement. As shown in Table 2, we integrated the diffusion augmentation method into the original train (no augmentation data) without any self-refinement, resulting in a 0.2% decrease in mIoU on the validation set. We speculate that this decline is due to the randomness of the diffusion process, which may have introduced low-quality augmented images. After incorporating image self-refinement during the generation phase of diffusion model, the performance improved by 2.4%, shows that our approach effectively enhances image quality. Finally, we introduced prompt self-refinement to diversify the background of the diffusion images, which further increased performance by 0.6%. It proves the importance of diverse and reliable prompts for image generation.
@@ 5 Conclusion
In this work, we propose an IAA method for WSSS. Unlike previous approaches that solely rely on the original training data, we employ an augmentation agent combined with diffusion and LLM knowledge to generate additional training images. These generated images are then combined with the original training data to create the final dataset, increasing the overall training data size. Furthermore, we design a self-refinement prompt and a self-refinement image module to enhance image quality. Ultimately, our data augmentation method achieves SOTA results in WSSS.
ACKNOWLEDGMENT

### v4

@@ 1 $P\leftarrow\text{generate\_prompt}(Y)$
3 if $person"\in Y$ then
4    $M\leftarrow\text{detect\_map}(X_{in},\text{human\_pose})$
@@
6    $M\leftarrow\text{detect\_map}(X_{in},\text{canny\_edge})$
@@ 8 $score_{df}\leftarrow\text{classifier\_score}(X_{df})$
9 if $score_{df}>\text{high\_quality\_threshold}$ then
@@ 3.4 Final Training Dataset of WSSS
After selecting the high-quality generated training samples, the synthetic dataset $\mathcal{D}_{aug}$ and the original dataset $\mathcal{D}_{origin}$ are combined to form an extended dataset $\mathcal{D}_{final}$ for the training of WSSS. The final training dataset is represented as:
$\mathcal{D}_{final}=\mathcal{D}_{origin}\cup\mathcal{D}_{aug}.$ (5)
This extended dataset $\mathcal{D}_{final}$ serves as a comprehensive training set that not only includes the original data but also incorporates the augmented data generated by our proposed method. The combination of these two datasets is critical for enhancing the diversity and quality of the training data, which in turn improves the performance of the model.
The synthetic data in $\mathcal{D}_{aug}$ is generated through our IAA method, leveraging advanced augmentation techniques such as LLMs and diffusion models. By generating realistic and varied synthetic images, $\mathcal{D}_{aug}$ complements the original dataset $\mathcal{D}_{origin}$, providing a broader range of visual scenarios for the model to learn from. This is especially valuable for weakly supervised tasks where labeled data is scarce or difficult to obtain.
The integration of $\mathcal{D}_{aug}$ into the training process allows the model to better generalize to different environments and conditions, improving segmentation accuracy in more challenging scenarios. As a result, the final dataset $\mathcal{D}_{final}$ not only increases the quantity of the training data but also significantly enhances its variety, making it a crucial factor for improving the overall performance of WSSS.
Furthermore, this extended dataset $\mathcal{D}_{final}$ is used to train our segmentation model, allowing it to benefit from both the original data’s consistency and the diversity of augmented data. The diversity introduced by $\mathcal{D}_{aug}$ provides the model with new perspectives that are essential for improving segmentation performance, particularly when dealing with complex and varied visual inputs.
@@ 4 Experiments
In this section, we first present the details of the dataset, evaluation metrics, and implementation. Next, we compare our IAA with state-of-the-art methods on the PASCAL VOC 2012 Everingham et al. [2010] and MS COCO 2014 benchmarks Lin et al. [2014]. Finally, we conduct ablation studies to demonstrate the effectiveness of the proposed method.
4.1 Experimental Settings
Table 1: Semantic Segmentation Performance Comparison (mIoU) on PASCAL VOC 2012.
Model Pub. Backbone mIoU (%) MCTformer Xu et al. [2022] CVPR22 DeiT-S 61.7 SIPE Chen et al. [2022] CVPR22 ResNet50 58.6 ViT-PCM Dosovitskiy et al. [2020] ECCV22 ViT-B/16 69.3 TSCD Xu et al. [2023] AAAI23 MiT-B1 67.3 SAS Kim et al. [2023] AAAI23 ViT-B/16 69.5 FPR Chen et al. [2023] ICCV23 ResNet38 70.0 ToCo Ru et al. [2023] CVPR23 ViT-B 70.2 Zhang et al. Zhang et al. [2023a] TIP23 Res2Net-101 71.0 WaveCAM Xu et al. [2024] TMM24 ViT-B/16 70.1 SFC Zhao et al. [2024] AAAI24 ViT-B/16 71.2 IACD Wu et al. [2024a] ICASSP24 ViT-B/16 71.4 PGSD Hao et al. [2024] TCSVT24 ViT-B/16 68.7 IAA Ours ViT-B/16 72.3
Table 2: Semantic Segmentation Performance Comparison (mIoU) on MS COCO 2014.
Model Pub. Backbone mIoU (%) MCTformer Xu et al. [2022] CVPR22 Resnet38 42.0 ViT-PCM Dosovitskiy et al. [2020] ECCV22 ViT-B/16 45.0 SIPE Chen et al. [2022] CVPR22 Resnet38 43.6 TSCD Xu et al. [2023] AAAI23 MiT-B1 40.1 SAS Kim et al. [2023] AAAI23 ViT-B/16 44.5 FPR Chen et al. [2023] ICCV23 ResNet38 43.9 ToCo Ru et al. [2023] CVPR23 ViT-B 42.3 SFC Zhao et al. [2024] AAAI24 ViT-B/16 44.6 IACD Wu et al. [2024a] ICASSP24 ViT-B/16 44.3 PGSD Hao et al. [2024] TCSVT24 ViT-B/16 43.9 IAA Ours ViT-B/16 45.3
Dataset and Evaluated Metric. Our experiments are conducted on two widely used datasets: the PASCAL VOC 2012 dataset Everingham et al. [2010] and the MS COCO 2014 dataset Lin et al. [2014]. The PASCAL VOC 2012 dataset consists of $21$ categories, including both foreground and background objects, and it is commonly extended using the SBD dataset Hariharan et al. [2011], which provides additional pixel-wise annotations for semantic segmentation tasks. The MS COCO 2014 dataset, on the other hand, includes $81$ classes, offering a more diverse and complex set of categories for training. For the PASCAL VOC 2012 dataset, we use a total of 10,582 images with image-level annotations for training and a separate validation set of 1,449 images. In the case of the MS COCO 2014 dataset, approximately 82,000 images are utilized for training, and around 40,000 images are reserved for validation, with training images annotated at the image level.
To evaluate the performance of our methods, we adopt the widely used mean Intersection-over-Union (mIoU) metric. The mIoU provides a comprehensive measure of the model’s ability to segment images accurately, considering both the true positives, false positives
Implementation Details. Our IAA method integrates knowledge from pre-trained GPT-4o Hurst et al. [2024], Stable Diffusion Rombach et al. [2022], and ControlNet Zhang & Agrawala [2023]. We utilize ViT-B/16 as the ViT model. To facilitate the training of the patch-based image classifier, images are resized to 384×384 as Kolesnikov & Lampert [2016], and the 24×24 encoded patch features are preserved as input. The model is trained for up to 80 epochs with a batch size of 16, using $\epsilon=0.9$ as the high-quality threshold for both text and image. Our final training dataset serves as input for the WSSS framework, while keeping all other settings consistent with ViT-PCM Rossetti et al. [2022]. The experiments were conducted using two NVIDIA 4090 GPUs. Finally, we used the same verification tasks and settings as ViT-PCM Rossetti et al. [2022].
4.2 Final Segmentation Performance
@@
Comparison with State-of-the-arts. In this work, we evaluated the final segmentation performance of our proposed IAA method on the PASCAL VOC 2012 and MS COCO 2014 datasets. As presented in Table 1, following the approach of other comparative methods, we trained our IAA model using a training set consisting of 10,582 images with only image-level annotations. For validation, we used 1,449 images to assess the performance of our final semantic segmentation results, where our method outperformed existing approaches. For the MS COCO 2014 dataset, we employed a similar strategy, utilizing approximately 82k images for training and around 40k images for validation to evaluate the segmentation performance. As shown in Table 2, our method demonstrated superior results on this dataset as well.
Visualization of results. Our IAA method effectively leverages an augmentation agent that integrates the power of GPT-4o and diffusion models to enrich the training data for WSSS. By combining the generative capabilities of these advanced models, our approach is able to create diverse and realistic synthetic images that complement the original dataset. This augmentation process not only increases the size of the training data but also enhances its variety, providing the model with a broader range of visual scenarios to learn from, which is crucial for improving segmentation accuracy in complex tasks. Fig. 4 presents examples of the augmented data that are used as additional training inputs for the model. These synthetic images, generated by the augmentation agent, mirror the characteristics of the original dataset while introducing new variations that would be difficult or costly to obtain through manual annotation. This process is essential in overcoming the limitations posed by the scarcity of labeled data in many real-world applications of WSSS. Furthermore, Fig. 5 showcases several visualization examples of the segmentation outputs produced by our method. These examples highlight the effectiveness of our approach in generating precise and accurate segmentation maps. It can be observed that our method consistently delivers segmentation results that are more refined and detailed compared to other approaches. The enhanced quality of the segmentation outputs demonstrates the ability of our augmented training data and self-refining modules to significantly improve the model’s performance, particularly in scenarios where weak supervision and limited labeled data are prevalent. Our results provide clear evidence that integrating advanced augmentation techniques with self-refinement mechanisms leads to a substantial improvement in segmentation accuracy and robustness.
Table 3: Ablation Study on Different Data Augmentation Models: Direct Stable Diffusion vs. Controlled Diffusion with Image Self-Refinement on PASCAL VOC 2012.
Original Training Data Direct Stable Diffusion Image Self-Refine (Controlled) on Val 69.3 69.0 72.3
@@ 4.3 Ablation Studies
Table 4: Ablation studies on main components of the proposed framework on the Pascal VOC 2012 val.
@@
Figure 6: Qualitative comparison of augmented images. Left: Direct stable diffusion result without structural guidance. Right: Image generated using ControlNet with Image Self-Refinement. The controlled method preserves the structure and introduces plausible variations, whereas the uncontrolled result deviates significantly from the original image.
We conduct comprehensive ablation studies to evaluate the individual and collective contributions of our proposed components: Prompt Self-Refinement (PSR) and Image Self-Refinement (ISR). As summarized in Tab. 4, we first integrate Diffusion Augmentation (DA) into the baseline model (original training without augmentation), which results in a marginal performance drop of 0.2% in mIoU on the validation set. This degradation can be attributed to the inherent randomness of the diffusion process, potentially introducing low-quality or semantically inconsistent augmented images. To address this limitation, we incorporate Image Self-Refinement (ISR) during the diffusion generation phase, which significantly improves the mIoU by 2.4%. This substantial gain demonstrates the effectiveness of ISR in enhancing the quality and semantic consistency of the augmented images. Furthermore, the addition of Prompt Self-Refinement (PSR) to diversify the background context of the generated images yields an additional performance boost of 0.6%, highlighting the importance of diverse and semantically meaningful prompts in guiding the diffusion process. The progressive performance improvement from 69.3% to 72.3% validates the synergistic effect of our proposed components, with ISR playing a particularly crucial role in mitigating the quality issues introduced by the raw diffusion process. These results underscore the significance of both image-level and prompt-level refinement in achieving robust and semantically consistent data augmentation for segmentation tasks.
Detailed Analysis on Image Self-Refinement
In addition, as shown in Table 3, we performed an ablation study comparing standard stable diffusion and our Controlled Self-Refined Diffusion approach. When using unconditioned stable diffusion, the generated images often differ significantly from the originals, which can introduce noise into the learning process and hinder performance. Even when applying an image selection mechanism, the absence of structural constraints leads to excessive deviation, ultimately reducing segmentation accuracy. In contrast, our Controlled Self-Refined Diffusion strategy preserves the core structure of the input image while incorporating subtle variations, resulting in improved model training and better segmentation outcomes. To further illustrate the differences in augmented image quality, we provide a visual comparison in Fig. 6. The direct diffusion result (left) introduces significant distortion and semantic inconsistency, whereas our Controlled Self-Refined Diffusion result (right) preserves the original structure while introducing meaningful variations. This qualitative difference further supports the effectiveness of our refinement strategy in producing semantically aligned and structurally faithful augmented data.
Table 5: (Table 5 in revised version) Normalized CLIP similarity score $(1+\cos)/2$ between original and augmented images over 100 samples from the airplane class on PASCAL VOC.
Image class Image Size Augmentation Method Mean CLIP Similarity Airplane 100 Direct Stable Diffusion 0.750 Airplane 100 Controlled Self-Refined Diffusion 0.893
To quantitatively assess semantic consistency between original and augmented images, we compute CLIP-based cosine similarity scores over 100 randomly sampled airplane images. Specifically, we report the normalized similarity score as $(1+\cos(\theta))/2$, where $\cos(\theta)$ is the cosine similarity between the CLIP embeddings of the original and augmented images. As shown in Table 5, images generated using direct stable diffusion exhibit a lower average similarity of 0.750, indicating a larger deviation from the original image distribution. In contrast, images generated via our Controlled Self-Refined Diffusion approach achieve a much higher similarity of 0.893, demonstrating stronger semantic alignment. These results provide quantitative evidence that our strategy effectively reduces the domain gap and produces semantically consistent augmentations.
Table 6: Effectiveness of Self-Refinement on the airplane class. We report mIoU for three different prompt generation strategies to evaluate the impact of prompt quality and refinement.
Prompt Strategy Class mIoU
Fixed Prompt Baseline airplane 0.757
LLM Prompt w/o Self-Refine airplane 0.763
Ours: Self-Refined Prompt airplane 0.769
Detailed Analysis on Prompt Self-Refinement
To assess the necessity of the Self-Refinement module, we conduct an ablation study on the airplane class. As shown in Table 6, using a fixed prompt yields the lowest performance (0.757 mIoU), indicating that prompt diversity is crucial for effective image generation. When using LLM-generated prompts without refinement, we observe a small performance gain (0.763 mIoU), demonstrating the benefit of leveraging LLM capabilities.
Detailed Analysis on Inference Efficiency
Although our framework includes additional components during the training phase, it introduces no extra computation overhead during inference. As shown in Table 7, our method and the baseline both require approximately 0.95 seconds per image, totaling around 23 minutes for the entire PASCAL VOC 2012 validation set of 1449 images. This confirms that our design maintains deployment efficiency, making it suitable for real-world applications despite its training-time complexity.
Table 7: Comparison of inference time on the PASCAL VOC 2012 val set with and without IAA-enhanced images.
Setting Dataset Inference Time w/o IAA image PASCAL VOC 2012 val $\sim$23 minutes (0.95 sec/image) w/ IAA image PASCAL VOC 2012 val $\sim$23 minutes (0.95 sec/image)
Detailed Analysis on Quality Score Setting
To validate the choice of $\epsilon=0.9$, we conducted a human evaluation study across thresholds $\{0.7,0.8,0.9,0.95\}$. For each threshold, we randomly sampled 100 prompts and 100 images from three representative semantic categories (airplane, cat, and dog) in VOC and COCO. Human annotators were asked to assess whether each sample was semantically meaningful and relevant to its class. As shown in Table 8 and Table 9, the acceptance rate increases with the threshold and saturates at $\epsilon=0.9$, where all samples were judged as high-quality. Further increasing $\epsilon$ to 0.95 offers minimal perceptual improvement while significantly increasing generation cost. This confirms that $\epsilon=0.9$ achieves a good trade-off between quality and efficiency.
Table 8: Prompt quality at different $\epsilon$ thresholds (evaluated on 100 randomly sampled prompts per threshold on PASCAL VOC).
Category $\epsilon=0.7$ $\epsilon=0.8$ $\epsilon=0.9$ $\epsilon=0.95$ Airplane 74% 88% 100% 100% Cat 70% 85% 100% 100% Dog 76% 89% 100% 100% Average 73.3% 87.3% 100% 100%
Table 9: Image quality at different $\epsilon$ thresholds (evaluated on 100 randomly sampled images per threshold on PASCAL VOC).
Category $\epsilon=0.7$ $\epsilon=0.8$ $\epsilon=0.9$ $\epsilon=0.95$ Airplane 68% 89% 100% 100% Cat 71% 86% 100% 100% Dog 69% 90% 100% 100% Average 69.3% 88.3% 100% 100%
@@ 5 Conclusion
In this work, we propose the IAA method for WSSS, addressing key challenges through innovative data augmentation. Unlike traditional methods relying solely on original training data, our approach introduces an augmentation agent that leverages diffusion models and LLMs to generate diverse synthetic images consistent with the original dataset. These images enrich the training data, providing greater variety and enabling the model to learn from a broader range of visual scenarios. By seamlessly integrating augmented images with the original data, our method significantly expands the dataset without additional manual labeling, mitigating the issue of limited labeled data. This expansion not only improves model performance but also reduces overfitting by exposing the model to more diverse and realistic contexts. Additionally, we design two key modules to enhance image quality and generalization: the self-refinement prompt and the self-refinement image module. The self-refinement prompt iteratively improves generated background prompts, ensuring task relevance, while the self-refinement image module iteratively optimizes generated images to meet desired characteristics. Together, these modules enhance the data augmentation process, producing higher-quality and more reliable training samples for improved segmentation performance.
Ultimately, our data augmentation method demonstrates SOTA results in weakly supervised semantic segmentation, setting a new benchmark for the field. By combining cutting-edge techniques in data generation, prompt refinement, and image quality enhancement, we offer a robust solution for overcoming the challenges of limited labeled data, improving model performance, and achieving more accurate segmentation results. This approach opens up new avenues for future research in WSSS, particularly in expanding the capabilities of weakly supervised models to handle a wider variety of real-world scenarios.
In future work, we plan to extend our evaluation to additional datasets beyond PASCAL VOC and COCO, especially those involving more diverse, complex, and real-world scenes. This will help assess the robustness and generalization capacity of our method. Furthermore, we aim to explore the adaptation of our augmentation framework to downstream tasks such as object detection and instance segmentation, and to investigate its effectiveness in low-resource or domain-shifted environments. These directions will provide deeper insights into the scalability and applicability of our approach in broader practical contexts.
6 Acknowledgements